OpenStudy (anonymous):

1. Joseph borrows $10,000 from the shady loan shark, Mr. Wildcat. Mr. Wildcat tells him he must repay the loan in 1 month and pay an additional $1000 in interest. If Mr. Wildcat is charging simple interest on his loan, what is the annual interest rate?

OpenStudy (anonymous):

$12,000

OpenStudy (anonymous):

the answer will be a % because the question asks for 'interest rate'

OpenStudy (anonymous):

I got -1.1 and im pretty sure that is wrong since its negative

OpenStudy (mathstudent55):

$1,000 is what percent of $10,000? \( percent = \dfrac{part ~amount}{total ~amount} \times 100 \) First, use the formula above to find what percent of $10,000 is $1,000.

OpenStudy (anonymous):

10%?

OpenStudy (mathstudent55):

The part amount is $1,000. The total amount is $10,000. Correct.

OpenStudy (mathstudent55):

He is charging 10% interest in 1 month. 1 year = 12 months How much interest is he charging in 1 year?

OpenStudy (anonymous):

120?

OpenStudy (mathstudent55):

Correct. The annual interest rate is 120%.

OpenStudy (anonymous):

so thats the answer? wow that seems so easy!

OpenStudy (mathstudent55):

That's it.

OpenStudy (anonymous):

well thank you very much!

OpenStudy (mathstudent55):

You're very welcome.

OpenStudy (anonymous):

Why is it 120%? Can someone give me the formula to get that?

OpenStudy (anonymous):

@Mathstudent55

OpenStudy (mathstudent55):

The formula for percentage is shown above. I'll do it here again. \(percent = \dfrac{part}{total} \times 100 \%\) In this case, the total amount of the loan was $10,000. That is the "total" in the percent formula. The interest paid was $1,000. This is the "part" in the formula. We want to know $1,000 is what percent of $10,000. By using the formula, we get: \(percent = \dfrac{$1000}{$10,000} \times 100 \%\) \( percent = \dfrac{1}{10} \times 100 \%\) \(percent = 10 \%\) By using the formula, we now know that $1,000 is 10% of $10,000. That means that the loan had an interest rate of 10% in one month. Since there are 12 months in a year, and the problem is asking for the annual interest rate, we must multiply the monthly interest rate by 12 to get \(12 \times 10 \% = 120 \%\) The final answer is: The annual interest rate is 120%.